OpenStudy (kimjafo):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128

Directrix (directrix):

Have you found the common ratio yet?

OpenStudy (kimjafo):

No

Directrix (directrix):

What do you multiply -2 by to get 8?

Directrix (directrix):

@KimJaFo

OpenStudy (kimjafo):

8

OpenStudy (kimjafo):

I mean 4

Directrix (directrix):

-4, I think.

Directrix (directrix):

What do you multiply 8 by to get -32

OpenStudy (kimjafo):

-4

Directrix (directrix):

And, to get 128, you multiply -32 by -4. So, the common ratio seems to be -4

OpenStudy (kimjafo):

Yes

Directrix (directrix):

Do you have answer options? If so, we can test them against the terms of the sequence.

OpenStudy (kimjafo):

No, I don't sadly. Thats why im having trouble with it.

Directrix (directrix):

To get the signs to alternate, we will have to use (-1)^n where n is an integer >0. So, if n = 1, then (-1)^n will = -1. If n = 2, then (-1)^n will - + 1 That makes the signs alternate negative, positive, negative, positive and so on.

Directrix (directrix):

Now, for the value of the terms: -2, 8, -32, 128 Would you write them as powers of 2? Ignore the signs for now. For example, 2 = 2^1 and 8 = 2^3 You do 32 and then 128 as powers of 2, okay?

OpenStudy (kimjafo):

okay